OpenStudy (anonymous):

5p - 4 = 8p + 4 solve the equation

OpenStudy (anonymous):

add 4 to both sides. Then subtract 8p from both sides Finally divide both sides by -3

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

-6=p

OpenStudy (anonymous):

p=-8/3

OpenStudy (anonymous):

5p-4=8p+4 5p-8p=4+4 -3p=8 p=-8/3

OpenStudy (anonymous):

oh oops typo. in my question. supposed to be 5p - 14 = 8p + 4

OpenStudy (anonymous):

add 14 to both sides then subtract 8p from both sides finally divide both sides by -3

OpenStudy (anonymous):

you should get p=-6

OpenStudy (anonymous):

thanks, i got that right! as you can see above ^ haha before i realized the typo.